How many gallons of 80% antifreeze must be added to 90 gallons of 20% antifreeze to get a mixture of 70% antifreeze?
Answer by stanbon(75887)

How many gallons of 80% antifreeze must be added to 90 gallons of 20% antifreeze to get a mixture of 70% antifreeze?
------
alcohol + alcohol = alcohol
0.20*90 + 0.80x = 0.70(90+x)
---
Multiply thru by 100 to get:
20*90 + 80x = 70*90 + 70x
10x = 50*90
x = 450 gallong (amt. of 80% solution needed)
======================
Cheers,
Stan H.
======================